A parallel plate capacitor has square plates of side 5 cm and separated by a



distance of 1 mm. (a) Calculate the capacitance of this capacitor. (b) If a 10 V



battery is connected to the capacitor, what is the charge stored in any one of the



plates? (The value of εo = 8.85 x 10-12 Nm2 C-2)

1
Expert's answer
2022-04-28T16:42:00-0400

(a)

"C=\frac{\epsilon_0A}{d}=\frac{8.85*10^{-12}*(0.05)^2}{0.001}=2.2*10^{-11}\:\rm F"

(b)

"q=CV=2.2*10^{-12}*10=2.2*10^{-10}\rm C"